# Holistic Optimization Framework for FPGA Accelerators

### v3

@@ Abstract
This paper introduces Prometheus, a holistic framework that integrates task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution into a unified design space. Leveraging Non-Linear Problem methodologies, Prometheus explores this space to find solutions under resource constraints, enabling bitstream generation.
I Introduction
The emergence of customized accelerators has revolutionized modern computing by delivering significant energy efficiency and performance gains through specialization. Field-Programmable Gate Arrays (FPGAs) have gained prominence for their flexibility and high-performance potential. High-level synthesis (HLS) and source-to-source compilers [19, 81, 76, 32, 62, 50, 77, 31, 79, 38, 80, 43] have made hardware design more accessible by converting high-level code (e.g., C++, Python) with hardware-specific directives into hardware descriptions. However, their effectiveness heavily relies on the quality of the input code, the applied hardware directives, and the use of advanced code transformations and directive strategies to optimize data communication and achieve high Quality of Result (QoR).
Challenges in Existing Approaches  Several works have contributed to enhancing performance through Design Space Exploration (DSE) [58, 59, 67, 66, 64, 24, 82, 66] and source-to-source compilers [38, 79, 78, 60, 71]. Despite these advancements, significant challenges remain, notably in supporting Code Transformation, which involves efficiently optimizing code to fully exploit FPGA potential; Overlapping Communication and Computation, aimed at maximizing concurrency to enhance performance; Concurrent Task Execution, which ensures simultaneous execution of tasks without resource conflicts; and Exploiting Intra-Task Parallelism, leveraging parallel execution within tasks to further boost performance. Previous works address specific optimization challenges but do not tackle all aspects simultaneously, underscoring the need for comprehensive, integrated solutions. Moreover, generating synthesizable code that can be converted into a bitstream without human intervention is a persistent challenge in FPGA design. This complexity is exacerbated by the intricate architecture of FPGAs, which often incorporates multiple Super Logic Regions (SLRs). In addition to adhering to the overall resource constraints of the entire FPGA board, it is crucial to meet specific constraints of each individual SLR while the following factors need to be considered to generate a high-quality HLS solution.
Shared Buffer  A classical approach to improving performance involves using shared buffers, which cache data on-chip to facilitate reuse by various computational tasks. This method increases parallelism and is particularly effective for computation-bound kernels. Studies have shown the effectiveness of shared buffers [66, 64, 79, 77, 38, 31, 60, 59, 58]. However, on-chip memory limitations often prevent caching all necessary data, and overlapping computation and communication remains constrained. While some compilers like Merlin [77] and PASTA [35] support double buffering for single arrays, Merlin’s application is limited and context-specific, and PASTA lacks parallelism.
Dataflow  On FPGAs, dataflow enables seamless data movement between off-chip and on-chip memory and computational tasks. This method supports the overlap of communication and computation, enabling concurrent execution of tasks and allowing dependent tasks to proceed once data is available via FIFO (First-In-First-Out). Dataflow is highly effective for memory-bound kernels but limits parallelism due to the absence of shared memory among tasks. Each task is compiled independently, which simplifies the compilation process and can enhance operating frequency [26, 27, 28].
Integrating shared buffers with dataflow methods combines their strengths, boosting performance and reducing resource use. A specialized design space for automated bitstream synthesis streamlines the process by reducing manual rewriting of non-synthesizable code, requiring efficient global and per-task resource management.
Prometheus Approach  To address these limitations and advance high-level source-to-source compilation, we propose Prometheus. This framework integrates the benefits of shared buffers and dataflow, aiming to improve performance while managing resources efficiently to ensure bitstream generation. The main contributions are:
  • The introduction of a unified design space facilitates the development of a tiled dataflow architecture. This space incorporates task fusion, tiling strategies, loop permutation, overlapping of computation and communication automatically, and concurrent task execution.
  • Employing Non-Linear Problem methodology to navigate and discover theoretically optimal solutions within stringent resource constraints for generating bitstreams.
  • Automating our framework that is end-to-end comprehensive. It automatically generates a host file in OpenCL to verify correctness, runs Vitis HLS for report generation and RTL simulations, and generates the bitstream.
  • Conducting evaluations utilizing RTL simulation and on-board FPGA runs, delivering superior Quality of Results (QoR) compared to AutoDSE [67], Sisyphus [60], ScaleHLS [79], and Allo [14]. Importantly, it achieves these results without requiring DSE for HLS design generation.
II Background and Motivation
II-A HLS optimization
In High-Level Synthesis (HLS), directives such as unroll, pipeline, and array_partitioning pragmas play a crucial role. The unroll pragma enables parallel execution by duplicating statements, while pipeline stages loops for overlapping operations based on the initiation interval (II). array_partitioning divides arrays into on-chip memory banks, facilitating simultaneous data access and enhancing unrolling. Additional optimizations include shared buffers for reducing off-chip transfers, coarse-grained parallelism through duplicated processing elements, and ping-pong buffering to overlap data transfer and computation [18]. On AMD/Xilinx FPGAs, data transfers use bit widths up to 512 bits, with the maximum bit width determined by problem dimensions. For instance, transferring 216 floats with a 256-bit width (8 floats per transfer) requires 27 transfers. The dataflow pragma organizes computation into actors that communicate via FIFO queues, where each actor processes inputs and produces outputs as they become available [41].
Performance evaluation methods include estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation. Estimation and simulation avoid place-and-route but utilize board resources, while on-board evaluation imposes stricter constraints for bitstream generation.
II-B Limitation of existing methods
To demonstrate the limitations of the existing framework and methodology we use 3mm. As illustrated in Listing 1, two independent matrix multiplications compute matrices that are then utilized in the final matrix multiplication.
[…]

### v4

@@ Abstract.
This paper presents Prometheus, a holistic optimization framework that integrates key optimizations—including task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution—into a unified design space. By leveraging Non-Linear Programming (NLP) methodologies, Prometheus explores the optimization space under strict resource constraints, enabling automatic bitstream generation. Unlike existing frameworks, Prometheus considers interdependent transformations and dynamically balances computation and memory access.
We evaluate Prometheus across multiple benchmarks, demonstrating its ability to maximize parallelism, minimize execution stalls, and optimize data movement. The results showcase its superior performance compared to state-of-the-art FPGA optimization frameworks, highlighting its effectiveness in delivering high QoR while reducing manual tuning efforts.
High-Level Synthesis, NonLinear Programming, Compiler
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
1. Introduction
The rise of customized accelerators has revolutionized modern computing, driving significant improvements in energy efficiency and performance through hardware specialization. Among these accelerators, Field-Programmable Gate Arrays (FPGAs) have gained widespread adoption due to their flexibility, high performance, and adaptability across diverse domains, including machine learning, scientific computing, financial modeling, and embedded systems. Unlike fixed-function hardware such as Application-Specific Integrated Circuits (ASICs), FPGAs offer reconfigurability, enabling hardware adaptation to workload-specific requirements. However, achieving a good Quality of Result (QoR) on FPGAs remains a complex and resource-intensive process, requiring careful tuning of computation, memory access, and parallelism strategies.
To address this complexity, High-Level Synthesis (HLS) tools and source-to-source compilers (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Pouget et al., 2025b; Basalama and Cong, 2025) have emerged as critical enablers of FPGA adoption, allowing developers to write hardware-accelerated programs in high-level languages such as C++ and Python. These tools generate synthesizable hardware descriptions, leveraging pragmas (hardware directives) and code transformations to optimize execution. Despite their advantages, HLS-generated designs still require substantial manual tuning to achieve high QoR, as the compiler’s ability to optimize performance remains highly dependent on the structure and directives applied to the input code.
1.1. Challenges in FPGA Optimization
Despite significant advancements in HLS automation and FPGA design methodologies, achieving high-performance and resource-efficient FPGA implementations remains a challenging problem. This complexity arises from the intricate interactions between computation, memory access, and parallel execution strategies. Several key challenges must be addressed to unlock the full potential of FPGA acceleration.
Code Transformation and Design Space Exploration: FPGA performance is highly dependent on how computations are structured and scheduled. Optimizations such as loop tiling, loop permutation, and loop unrolling play a crucial role in exposing parallelism and improving memory access efficiency.
Task Concurrency and Resource Management: Maximizing throughput in FPGA designs requires efficient task concurrency, where multiple computation kernels execute in parallel. However, coordinating parallel tasks introduces challenges in managing resource contention, memory bandwidth, and routing complexity. Overlapping computations effectively demands intelligent scheduling strategies that minimize stall cycles while avoiding bottlenecks caused by limited on-chip storage and off-chip memory latency.
Computation-Communication Overlap: FPGA acceleration is often hindered by inefficient data movement between off-chip memory (DDR/HBM) and on-chip compute units. Many workloads experience severe performance degradation due to memory access bottlenecks rather than computational limitations. Dataflow architectures aim to address this by implementing FIFO-based streaming to pipeline data transfers, but they often fail to exploit intra-task parallelism efficiently. In contrast, shared buffer models (Pouget et al., 2025b; Ye et al., 2022) improve data reuse, but their rigid memory management strategies make it difficult to overlap computation and communication dynamically. A more adaptive approach is needed to orchestrate data movement in parallel with execution, reducing overall memory stalls.
Scalability and Multi-SLR Utilization: FPGA architectures, especially those using Stacked Silicon Interconnect (SSI) technology, feature multiple Super Logic Regions (SLRs) to enhance scalability. However, most frameworks, including Merlin-based tools (e.g., AutoDSE, HARP, NLP-DSE) and Sisyphus, restrict designs to a single SLR, leading to underutilization of FPGA resources. While designing across multiple SLRs is possible, it significantly increases routing congestion, timing closure failures, and bitstream generation complexity. Without an SLR-aware optimization strategy, large-scale tasks face a high probability of bitstream generation failures or degraded performance due to excessive inter-SLR communication overhead.
Comprehensive Design Space Exploration: Current design space exploration (DSE) methodologies are often narrow in scope, considering only a limited subset of optimizations at a time. Each transformation—whether it involves loop reordering, tiling, pipelining, or data partitioning—has a profound impact on the overall performance and resource usage. Frameworks that optimize only one aspect of the design fail to account for the interdependencies between different optimizations, leading to suboptimal results. To achieve maximum efficiency, FPGA designs require a global optimization approach that integrates all possible transformations while considering the constraints imposed by memory hierarchy, computation parallelism, and hardware utilization.
Addressing these challenges requires a unified optimization strategy that seamlessly integrates code transformation, task concurrency management, computation-communication overlap, and hardware-aware scheduling into a single framework. The proposed work aims to develop such an approach, ensuring high-performance FPGA implementations while maintaining synthesis feasibility.
1.2. Contributions and Prometheus Method
This work introduces Prometheus, a unified framework that automates FPGA design space exploration by integrating multiple optimizations into a single methodology. Unlike existing tools that optimize isolated aspects of FPGA acceleration, Prometheus holistically addresses code transformation, memory management, task concurrency, and hardware-aware scheduling. It leverages a hybrid execution model that balances shared memory reuse and dataflow streaming to maximize parallelism while minimizing memory overhead.
To explore the vast design space efficiently, Prometheus formulates the optimization process as a Non-Linear Programming (NLP) problem, enabling automatic selection of loop transformations, tiling strategies, pragma configurations, and memory partitioning policies. Unlike heuristic-based approaches, this formulation captures complex optimization interdependencies and ensures globally efficient designs. Additionally, Prometheus incorporates SLR-aware task scheduling, which partitions tasks across multiple Super Logic Regions (SLRs). This addresses a major limitation in prior works that restrict execution to a single SLR, leading to underutilization of FPGA resources. By integrating SLR-aware optimizations, Prometheus reduces routing congestion, improves scalability, and ensures successful bitstream generation.
The methodology follows a structured process: First, Prometheus performs affine code analysis to identify parallelism opportunities and dependencies. It then generates a dataflow graph and fuses tasks that share the same outputs to minimize unnecessary communication overhead. Next, Prometheus constructs a comprehensive design space that includes all key optimizations outlined in Section 2, such as task concurrency via dataflow, computation-communication overlap, and adaptive parallelism, while ensuring resource constraints are met for each SLR. The NLP-based design space exploration model, as described in Section 4, is then employed to identify the theoretically optimal set of transformations. Once the best configuration is identified, Prometheus automatically generates the optimized design. Finally, the system produces an OpenCL host code and compiles the FPGA bitstream with the HLS compiler.
The key contributions of this work are:
  • Unified FPGA optimization framework that jointly optimizes loop transformations, pragma insertion, task concurency and computation-communication overlap while considering interdependencies between computation and data movement.
  • Hybrid execution model that dynamically selects between shared buffering and dataflow streaming to maximize parallelism and efficient memory access.
  • NLP-based design space exploration that automates the selection of loop tiling, scheduling, and memory strategies for globally theoretical optimal performance.
  • SLR-aware scheduling and multi-SLR partitioning to balance routing complexity and FPGA resource utilization, overcoming single-SLR limitations in prior works.
  • End-to-end compilation and automation that generates optimized HLS-C++ code, OpenCL host code, and FPGA bitstreams with minimal manual intervention.
  • Comprehensive performance evaluation, demonstrating superior QoR compared to AutoDSE (Sohrabizadeh et al., 2021), Sisyphus (Pouget et al., 2025b), ScaleHLS (Ye et al., 2022), Stream-HLS (Basalama and Cong, 2025) and Allo (Chen et al., 2024).
2. Background and Motivation
Optimizing latency and performance in FPGA designs relies on a combination of code transformations and strategic insertion of directives within High-Level Synthesis (HLS) (Cong et al., 2018). This section explores various HLS optimization techniques, their role in enhancing parallelism and resource utilization, as well as their limitations and challenges.
Table 1 presents a comparative analysis of various frameworks built on top of AMD Vitis HLS. These frameworks are categorized based on their underlying methodologies, including Model-Free, AI-Based, Cost Model Communication, Cost Model Computation, and NLP-Based approaches. The table evaluates their capabilities across key optimization techniques and objectives.
Model Free AI Model Cost Model Communication Cost Model Computation NLP-Based
AutoDSE (Sohrabizadeh et al., 2021) HARP (Sohrabizadeh et al., 2023) PolyOpt-HLS (Pouchet et al., 2013a) ScaleHLS (Ye et al., 2022) / POM (Zhang et al., 2024) HeteroCL (Lai et al., 2019) / Allo (Chen et al., 2024) StreamHLS (Basalama and Cong, 2025) NLP-DSE (Pouget et al., 2025a) Sisyphus (Pouget et al., 2025b) Prometheus
Pragma Insertion Limit.
Code Transformation (Tiling) Limit Limit
Code Transformation (Loop Permutation) Limit Limit
Code transformation + pragma insertion (unified)
Task Concurrency Limit.
Dataflow
Computation-Communication Overlap Limit.
Data Packing
Padding (Communication)
Padding (Computation)
Hardware Aware Limit.
Hardware-feasible Limit. Limit. Limit. Limit.
Management of Off-Chip and On-Chip Memory Transfers
Objective Comm + Comp Comm + Comp Comm Comp - Comp Comm + Comp Comm + Comp Comm + Comp
Enumeration (AI, heuristics, …) Manual
Table 1. Comparison of HLS-Based FPGA Optimization Frameworks. This table summarizes key features supported by various frameworks, including pragma insertion, code transformations, task concurrency, dataflow modeling, and hardware-awareness, categorized by their underlying optimization strategies.
2.1. HLS Optimization
2.1.1. Pragma Insertion
HLS optimizations rely extensively on a range of directives that control loop execution, data access patterns, and computation scheduling to maximize performance. Among these, three fundamental pragmas—unroll, pipeline, and array partitioning—are particularly critical for improving parallelism and reducing execution latency.
The unroll directive enables the expansion of loop iterations in the hardware design, allowing multiple iterations to execute in parallel. By eliminating loop control overhead and enabling greater resource utilization, unrolling can significantly boost throughput. However, its effectiveness is limited by FPGA resource constraints, such as available DSP slices, BRAM, and routing congestion.
The pipeline pragma restructures loop execution to allow overlapping operations, thereby reducing the initiation interval (II)—the number of cycles required to launch consecutive iterations. This directive is essential for achieving high throughput in applications with iterative computations, such as matrix multiplications and stencil operations. Careful tuning of pipeline depth and initiation interval is necessary to balance resource usage while avoiding performance bottlenecks due to memory access contention.
Array partitioning is another crucial optimization that enhances memory access parallelism. By splitting large arrays into smaller banks stored in separate on-chip BRAM blocks or distributed memory structures, this pragma allows simultaneous data accesses, facilitating more effective unrolling and pipelining. Without partitioning, memory conflicts and bandwidth limitations can restrict performance, especially with multiple concurrent memory accesses.
Beyond these fundamental pragmas, additional directives such as loop flattening, and dependency pragmas can further optimize execution. Loop flattening combines nested loops into a single loop to improve hardware efficiency. Dependency pragmas help manage data dependencies to ensure efficient scheduling without unnecessary synchronization delays.
Pragma insertion has been extensively studied using various methodologies to optimize performance and resource utilization in HLS designs. These approaches include bottleneck analysis, as employed by AutoDSE (Sohrabizadeh et al., 2021), performance estimation through Graph Neural Networks (GNN) in frameworks like Harp (Sohrabizadeh et al., 2023), and cost model optimization leveraging Non-Linear Programming (NLP) in NLP-DSE (Pouget et al., 2025a, 2024).
Each method presents distinct advantages and limitations. AutoDSE achieves full accuracy by running the HLS compiler for each configuration, ensuring precise performance measurements. However, this approach is computationally expensive and time-consuming, making it impractical for rapid design exploration. HARP, on the other hand, can estimate performance across configurations without direct compilation, significantly reducing evaluation time. However, it requires extensive training datasets and fine-tuning for each new kernel to accurately model compiler behavior. While this approach enables learning-based adaptation to different designs, its effectiveness depends heavily on data quality and model calibration. Conversely, NLP-DSE explores the entire design space efficiently by formulating the optimization as a Non-Linear Programming problem. This method allows for rapid exploration and selection of theoretical optimal configurations. However, it relies on a theoretical cost model, which may introduce inaccuracies if the compiler behaves unpredictably or if certain optimizations are not accurately captured by the model. Consequently, while NLP-DSE provides a fast and scalable solution, its reliability depends on the fidelity of the cost model to actual HLS compilation behavior.
While pragma insertion is a powerful optimization technique, its effectiveness is inherently dependent on the interaction between memory hierarchy, computational resources, and FPGA-specific constraints. Moreover, its impact is tightly coupled with the program’s execution schedule, meaning that without appropriate code transformations, the benefits of pragma insertion remain limited. Effective optimization requires a synergy between pragma directives and structural modifications to the code to fully exploit FPGA parallelism and resource utilization.
2.1.2. Code Transformation
Code transformations are fundamental for optimizing execution on FPGAs. While pragma insertion plays a crucial role in enhancing performance, it must be complemented by effective code transformations to fully exploit FPGA resources. Techniques such as loop reordering (permutation), task fusion, and data tiling enhance data locality and increase parallelism, thereby improving both computational efficiency and memory access patterns.
Various code transformation strategies have been developed specifically for FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). While transformations originally designed for CPUs and GPUs achieve substantial performance gains by optimizing for their respective architectures, they do not inherently align with FPGA requirements, which prioritize fine-grained parallelism and efficient resource utilization. Several studies have utilized Pluto (Bondhugula et al., 2008), a leading compiler framework originally designed for CPU optimizations, to transform FPGA kernels (Zhao et al., 2022; Zhao and Cheng, 2021). While Pluto excels at tiling and minimizing dependencies to enhance memory reuse, its direct application to FPGA optimization is constrained due to the fundamental differences in optimization strategies required for CPUs and FPGAs. Unlike CPUs, where memory hierarchy and cache locality are primary concerns, FPGA optimization focuses on maximizing parallelism, minimizing resource contention, and efficiently utilizing on-chip memory, making Pluto’s conventional transformation techniques less effective in this context. Conversely, studies such as (Pouchet et al., 2013b; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on code transformations tailored to specific FPGA performance goals. The work in (Pouchet et al., 2013b) aims to reduce communication overhead between off-chip and on-chip memory, achieving superior Quality of Results (QoR) for memory-bound kernels. Meanwhile, research efforts in (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on optimizing pipelining strategies to maximize instruction-level parallelism and resource utilization. Sisyphus (Pouget et al., 2025b) introduces a unified approach by integrating code transformation and pragma insertion into a single optimization problem. By formulating this as a Non-Linear Programming (NLP) problem, Sisyphus efficiently explores the design space to identify theoretical optimal configurations, streamlining FPGA acceleration while maintaining a balance between computation and memory access efficiency.
2.1.3. Task Concurrency
HLS tools like Vitis HLS support the dataflow pragma, which structures computations into actors that communicate through FIFO queues. This approach allows overlapping execution of multiple tasks, significantly reducing overall latency . By enforcing a producer-consumer model, dataflow scheduling enables each task to process data as soon as it becomes available, rather than waiting for an entire stage to complete, which is particularly beneficial.
For computational kernels such as 3mm (Listing 3), the dataflow paradigm allows the first two matrix multiplications to execute in parallel while the third begins as soon as its required inputs are produced. This overlapping of execution helps maximize throughput and minimize idle time for computing units. Additionally, dataflow optimizations facilitate task-level parallelism, allowing independent tasks to run concurrently across multiple compute resources, such as DSP blocks and BRAM, ensuring efficient utilization of available FPGA resources.
However, despite its advantages, pure dataflow parallelism presents several challenges. One major limitation is intra-task parallelism, which is constrained by the reliance on FIFO-based communication. Since each FIFO can only transfer up to 512 bits per cycle (as this is the maximum off-chip memory bitwidth supported), this inherently restricts the amount of data that can be processed concurrently within a task. To further enhance intra-task parallelism, alternative approaches must be explored.
Stream-HLS (Basalama and Cong, 2025) attempts to address this limitation by increasing the number of FIFOs connecting two tasks, assuming that all data resides on-chip. While this approach simplifies certain aspects of the design space, it is neither scalable nor generalizable for transferring data from off-chip to on-chip memory. By using $n$ FIFOs to transfer $n$ data elements in parallel between two tasks, the method significantly increases resource consumption without proportionally improving efficiency. The additional FIFOs complicate the design, potentially leading to routing congestion and excessive hardware overhead.
Moreover, modern FPGAs feature up to 32 off-chip memory banks, making the multi-FIFO approach impractical for handling off-chip data transfers. A more effective and scalable strategy must be developed to optimize intra-task parallelism while ensuring efficient communication between off-chip and on-chip memory, without introducing unnecessary complexity or resource constraints.
2.1.4. Shared Buffering
Shared buffering is a critical technique in FPGA memory optimization, enabling efficient data access and reuse across multiple computational units. It can be employed at a global level, as seen in frameworks like AutoDSE (Sohrabizadeh et al., 2021), Sisyphus (Pouget et al., 2025b), and ScaleHLS (Ye et al., 2022), or within individual dataflow tasks to enhance execution efficiency.
This approach involves preloading data buffers into on-chip memory, allowing multiple computations to access shared data without redundant transfers. By reducing memory access latency and improving bandwidth utilization, shared buffering facilitates high parallelism and optimizes overall performance. It is particularly beneficial in applications which are computation bound.
However, shared buffering presents challenges in maintaining concurrency. While it enables efficient data reuse, it can introduce synchronization overhead, especially when multiple tasks attempt to access the same memory region simultaneously. Managing concurrent access requires arbitration mechanisms, which can lead to increased latency and potential bottlenecks. Additionally, routing congestion can occur due to high interconnect demands, limiting scalability in complex FPGA designs.
One limitation of shared buffering is its impact on initiation interval (II) in pipelined architectures. Unlike dataflow-based designs that rely on FIFO queues for seamless data streaming, shared buffering requires explicit read and write coordination, which may introduce stalls if not carefully managed. Moreover, FPGA resource constraints, such as limited BRAM and URAM availability, impose restrictions on buffer size and allocation strategies.
To address these challenges, advanced techniques such as double-buffering, memory partitioning, and adaptive scheduling have been explored. Double-buffering enables overlapping computation with data transfer, reducing idle cycles. Memory partitioning distributes data across multiple banks to alleviate contention, while adaptive scheduling dynamically assigns buffer access based on task priority and workload demands.
By effectively integrating shared buffering with intelligent memory management strategies, FPGA designs can achieve a balance between computational parallelism and efficient memory access. Future advancements should focus on automated buffer allocation techniques and dynamic access pattern optimization to further enhance performance in high-level synthesis (HLS) workflows.
2.1.5. Computation-Communication Overlap
Overlapping computation and communication is crucial for high-performance FPGA designs. Techniques such as double buffering (ping-pong buffering) and advanced data tiling help mask communication latency while keeping computational units busy. Managing data transfers efficiently between on-chip and off-chip memory ensures that processing units remain active without waiting for data.
2.1.6. Data Packing and Padding
Modern FPGA architectures support high-bandwidth data transfers (up to 512-bit wide on AMD/Xilinx FPGAs). Data packing and padding optimize memory alignment, reducing the number of required memory cycles. For instance, transferring 216 floating-point values using a 256-bit width (8 floats per transfer) requires 27 cycles, making efficient packing crucial for minimizing overhead.
To fully exploit data packing, padding must be considered to enable even faster data transfers. Additionally, padding is valuable for achieving finer control over parallelism and resource utilization by expanding the available design space for loop unrolling, thereby enhancing computational throughput and efficiency.
Padding for Communication
Padding must be considered to increase the flexibility of data packing. The original size of the data may impose restrictions on packing efficiency, but by introducing padding, we create a larger space that may allow for a more efficient transfer. However, padding is not a free optimization, as increasing it also increases the amount of data that needs to be transferred.
For example, in the code presented in Listings 1 and 2, where $J=190$, transferring all elements along the second dimension (iterated by loop $j$) of array $B$ onto the chip is constrained by the available memory bandwidth. Without padding, the maximum transfer rate is $64$ bits per cycle, as $190\times 32$ is divisible by $64$ but not by $128$. However, by introducing padding with $P=2$ and adjusting $J$ to $192$ ($190+2$), the data alignment enables a significantly higher transfer rate of $512$ bits per cycle, as $192\times 32$ is now perfectly divisible by $512$. This optimization maximizes memory bandwidth utilization, reducing transfer latency and improving overall data throughput.
Padding for Computation
Padding can also be used for computation, if we take a similar method than Sisyphus (Pouget et al., 2025b) which tiled and unroll the intra-tile loops which correspond to the transformation from Listing 1 to Listing 2. We will have an unroll factor which correspond to $I1\times J1\times K1$ but we want to have $I1$ divide $I$, $J1$ divide $J$ and $K1$ divide K because we do not want to use extra ressource to only compute a partial tile which would correspond to the execution of the last tile which is not complete. For example if we have a trip count of $190$ and we use an unroll factor of $8$ then we will execute $184\times 8$ a full tile and then 6 iterations for the partial tile.
But we restrict the unroll factor to be divisible of the trip count we restrict considerably the space of possibilities. So the solution is to pad which allow to considerably increase the space of possibilities. Hence with a trip count of $190$ the possible unroll factor are include in $UF\in\{1,2,5,10,19,38,95,190\}$ but with padding of 2 (a trip count of 192) the space become $UF\in\{1,2,3,4,6,8,12,16,24,32,48,64,96,192\}$. And of course all pading can be considered. Padding enables finer control over unrolling and resource utilization, allowing for more flexible and efficient hardware implementation while optimizing computational parallelism.
1for (i = 0; i < I; i++)
2 for (j = 0; j < J; j++)
3 for (k = 0; k < K; k++)
4 C[i][j] += A[i][k] * B[k][j];
Listing 1: Baseline Implementation of Matrix Multiplication in C. This code depicts a naive triple-nested loop structure used for matrix-matrix multiplication, serving as the unoptimized reference for further transformations.
1for (i0 = 0; i0 < I0; i0++)
2 for (j0 = 0; j0 < J0; j0++)
3 for (k0 = 0; k0 < K0; k0++)
4 for (i1 = 0; i1 < I1; i1++)
5#pragma HLS unroll
6 for (j1 = 0; j1 < J1; j1++)
7#pragma HLS unroll
8 for (k1 = 0; k1 < K1; k1++)
9#pragma HLS unroll
10 C[i0*I1 + i1][j0*J1 + j1] += A[i0*I1 + i1][k0*K1 + k1] * B[k0*K1 + k1][j0*J1 + j1];
Listing 2: Matrix Multiplication with Loop Tiling and Fully Unrolled Intra-Tile Computation. This implementation showcases a performance-optimized matrix multiplication using loop tiling to enhance data locality and fully unrolled intra-tile loops to expose fine-grained parallelism.
2.2. Performance and Hardware Considerations
2.2.1. Performance Evaluation
Performance evaluation in HLS is a multi-stage process that assesses different aspects of a design’s efficiency and feasibility. It is typically conducted at three levels: estimation (e.g., Vitis HLS reports), RTL simulation, and FPGA-based evaluation. Each of these stages provides critical insights, but they vary in accuracy, speed, and resource requirements.
  1. (1)
    Estimation (HLS Reports): The first level of evaluation relies on estimation tools provided by HLS compilers such as Vitis HLS. These reports generate quick approximations of key performance metrics, including latency, resource utilization (DSPs, LUTs, BRAMs), and throughput. While these estimations are useful for early-stage design exploration, they do not account for low-level placement and routing effects, which can significantly impact real-world performance. The optimistic assumptions of HLS reports often fail to capture routing congestion, memory access delays, and other architectural bottlenecks.
  2. (2)
    RTL Simulation: To achieve more accurate performance insights, RTL simulation is performed after HLS compilation. This stage involves generating synthesizable RTL code, which is then simulated using hardware description language (HDL) tools. RTL simulation provides cycle-accurate performance metrics, allowing designers to analyze pipeline behavior, data dependencies, and execution timing. However, while RTL simulation models hardware more precisely than HLS reports, it does not incorporate real-world FPGA constraints such as clock tree synthesis, interconnect delays, and power distribution, which can affect the final implementation.
  3. (3)
    FPGA-Based Evaluation: The most accurate method of performance evaluation involves deploying the design onto an actual FPGA. This process includes place-and-route, bitstream generation, and execution on hardware. FPGA-based evaluation provides real execution metrics such as operating frequency, power consumption, and effective memory bandwidth. It also reveals practical challenges, including timing closure issues, placement inefficiencies, and resource contention that cannot be detected in earlier stages. This level of evaluation is essential for validating Quality of Results (QoR) and ensuring the design meets real-world constraints.
The transition from RTL simulation to FPGA implementation often introduces additional challenges, such as increased routing complexity, unexpected timing violations, and suboptimal resource utilization. These discrepancies arise due to the abstraction gap between HLS-generated code and final FPGA hardware constraints. To mitigate these issues, it is essential to develop robust tools that enable rapid design regeneration by modifying specific configurations within a well-defined portion of the design space. Such tools would enable engineers to efficiently iterate on design parameters without requiring a complete code regeneration, which can be time-consuming and may result in a significantly different design that still fails to generate a valid bitstream. Having the ability to selectively modify only the congested parts of the design while preserving the rest of the configuration would be highly valuable, ensuring faster convergence toward an optimized and feasible FPGA implementation.
2.2.2. Hardware Awareness and Resource Constraints
Many existing studies conclude their evaluations at the Vitis HLS report or RTL simulation stage, overlooking the critical impact of placement and routing constraints. However, as the design progresses toward bitstream generation, the available design space becomes increasingly constrained. Addressing hardware limitations early in the HLS process is essential to avoid costly design iterations and ensure convergence toward a feasible and deployable FPGA implementation.
Additionally, Super Logic Region (SLR)-aware optimizations are crucial for multi-SLR FPGA architectures. A Super Logic Region (SLR) is a physically distinct section of an FPGA die in Stacked Silicon Interconnect (SSI) technology. Each SLR contains a subset of the device’s computational resources, including DSP, LUT and FF, similar to those in monolithic FPGA devices.
Effectively mapping tasks across SLRs is critical for balancing resource utilization and minimizing routing congestion, as inter-SLR communication introduces additional latency and can become a performance bottleneck. By integrating SLR-aware task partitioning, optimized scheduling, and efficient data movement strategies, FPGA designs can achieve higher scalability, improved parallelism, and enhanced synthesis feasibility. Ensuring that computation and memory access patterns align with the physical structure of multi-SLR FPGAs is essential for achieving high QoR while maintaining timing closure and efficient resource allocation.
Some recent frameworks, such as RapidStream-TAPA (Guo et al., 2023a, b) and PASTA (Khatti et al., 2023), aim to improve the design scalability and performance of HLS programs on modern multi-die FPGAs through hardware-aware co-optimization of HLS and physical design. While RapidStream-TAPA focuses on latency-insensitive, FIFO-based communication and exploits coarse-grained floorplanning and pipelining for timing closure and parallel compilation, PASTA extends this model to include buffer-based inter-task communication using a generalized channel abstraction. Despite their advances, both frameworks still require a manually optimized task-parallel input code, which demands substantial expertise in HLS and hardware design, limiting their accessibility for non-expert users.
2.3. Challenges
While existing HLS optimizations provide substantial performance improvements, several challenges remain. One of the primary limitations is that many prior works restrict their optimization space to specific techniques or separate their optimization processes into multiple independent steps. This fragmented approach can lead to incoherent design decisions, where optimizations applied at one stage may not align with or may even counteract those applied at subsequent stages. Refer to Table 1 for a detailed comparison of these approaches.
A common oversimplification in many frameworks is the assumption that all data resides on-chip. While this simplifies memory access patterns in design space exploration, it often results in low Quality of Results (QoR) when deployed on real hardware. This assumption prevents frameworks from incorporating crucial tiling strategies, which are essential for optimizing data locality, reducing memory access overhead, and ensuring efficient off-chip communication. When off-chip memory management is treated as a separate optimization step, decisions made in earlier phases may restrict the effectiveness of later memory optimizations, leading to suboptimal performance and resource utilization.
The limitations of current HLS optimization frameworks that rely on shared buffering are exemplified through the 3mm kernel. It consists of two independent matrix multiplications, whose outputs serve as inputs for a final matrix multiplication.
As illustrated in Listing 3, each matrix multiplication consists of deeply nested loops that are well-suited to parallelization techniques such as loop unrolling, pipelining, and computation-communication overlap. However, conventional shared buffering strategies often fall short in effectively leveraging concurrency, primarily due to their limited exploitation of dataflow principles and insufficient overlap between computation and communication.
[…]